# Efficient Image Pre-Training with Siamese Cropped Masked Autoencoders

### v1

@@ Abstract
Self-supervised pre-training of image encoders is omnipresent in the literature, particularly following the introduction of Masked autoencoders (MAE). Current efforts attempt to learn object-centric representations from motion in videos. In particular, SiamMAE recently introduced a Siamese network, training a shared-weight encoder from two frames of a video with a high asymmetric masking ratio (95%). In this work, we propose CropMAE, an alternative approach to the Siamese pre-training introduced by SiamMAE. Our method specifically differs by exclusively considering pairs of cropped images sourced from the same image but cropped differently, deviating from the conventional pairs of frames extracted from a video. CropMAE therefore alleviates the need for video datasets, while maintaining competitive performances and drastically reducing pre-training time. Furthermore, we demonstrate that CropMAE learns similar object-centric representations without explicit motion, showing that current self-supervised learning methods do not learn objects from motion, but rather thanks to the Siamese architecture. Finally, CropMAE achieves the highest masking ratio to date (98.5%), enabling the reconstruction of images using only two visible patches. Our code is available at https://github.com/alexandre-eymael/CropMAE.
@@ 1 Introduction
Figure 1: CropMAE self-supervised learning pre-training. Given an input image ($V_{1}$), a second image is generated by performing a random crop and, optionally, a horizontal flip on the original image ($V_{2}$). We then patchify [14] both views and mask [33, 25] an extremely high portion of the second image (above 98.5%). Both views are encoded by a Siamese [5] ViT encoder, with added positional embedding [14]. A transformer [20] decoder reconstructs the masked image $R$ using self-attention layers on the tokens of the masked image and cross-attention layers between the tokens of the masked and unmasked images.
@@
Recently, Siamese Masked autoencoders (SiamMAE) [23] achieved state-of-the-art performance in numerous propagation tasks [56, 39, 30] by learning object-centric representations from videos. This method leverages a Siamese encoder [5] to process pairs of frames that are asymmetrically masked. Despite its impressive performance, SiamMAE faces two main limitations: firstly, it is designed to only process video frames, not standalone images. Yet, image datasets are typically orders of magnitude larger than video datasets, and less computationally expensive to decode, making image-based pre-training more effective and scalable than video-based pre-training. Secondly, while SiamMAE reduces the need for the intense data augmentation found in contrastive learning methods, it still requires learning a conceptual understanding of the visual world, similar to most MIM techniques, thus requiring extensive training ($2{,}000$ epochs) on large datasets such as K400 [32] to reach state-of-the-art performances.
@@
In this work, we propose a novel self-supervised learning method, called CropMAE, that reframes the siamese-based paradigm introduced in SiamMAE in order to alleviate the need for video dataset, while keeping competitive performances on downstream tasks. Specifically, we use random views of the same image to simulate viewpoint changes, object transformations and motion, and occlusions. Our method can therefore leverage both image and video datasets, and train at a significantly faster pace than SiamMAE. Moreover, we demonstrate that CropMAE learns meaningful object-centric representations for downstream video tasks without explicit motion. Finally, unlike most MIM techniques, the pretext task of CropMAE is directly tractable based on the visible frame without the need to learn conceptual information about the world, which we believe is the reason for its faster training. An overview of our method is presented in Figure 1.
@@
Contributions. We summarize our contributions as follows. (i) We introduce a novel pre-training method, CropMAE, based on sole images, which alleviates the need for video decoding and significantly accelerates training. The novel pretext task we introduce learns faster while quickly reaching good performances. (ii) We empirically demonstrate the feasibility of learning meaningful representations for downstream video tasks from still images or data distributions traditionally not associated with videos. Notably, this approach yields better results than training directly on video frames. (iii) We show, for the first time, that employing an extremely high masking ratio (98.5%, i.e., using only two visible patches for a ViT/16), surpassing those explored in existing studies, can be optimal and generate a meaningful and challenging self-supervised task.
@@ 2 Related Work
Visual representation learning. Visual self-supervised learning focuses on learning rich and generalizable representations of images or videos. This is typically achieved through pretext tasks [36, 37, 43, 8], enabling the learned representations to be applicable to a broad set of downstream tasks [15, 12, 56], either by fine-tuning the learned models for specific tasks, or by freezing the weights and training a linear classifier or an MLP on top of it. Key downstream tasks in the visual domain include image classification [8, 22, 6, 10, 52, 55, 3, 25, 19, 38], video classification [45, 48, 17, 19, 16, 38], object detection [22, 10, 25], and video segmentation [6, 10, 4, 23, 31]. Our method, CropMAE, is a new visual self-supervised representation learning method for propagation tasks [39, 56, 30].
@@
Contrastive Self-Supervised Learning. Contrastive self-supervised learning [24] has been recognized as an effective method for feature extraction, applicable both to images [10, 22] and videos [42, 11]. This approach encourages the encoder to learn robust representations of the input data by minimizing the distance between representations of different augmented versions of the same image. Initially, it was common to enforce distinct images to have different representations in order to avoid representation collapse [13, 50, 8]. However, subsequent discoveries [26, 22] have shown that robust learning can be achieved even without imposing this constraint. Contrastive self-supervised learning has also been widely used for correspondence learning [49, 29], as it inherently learns to build representations that are invariant and robust to perturbations. Contrary to contrastive learning, CropMAE does not extensively rely on data augmentations and is not subject to representation collapse issues.
@@
Masked Image Modeling. Drawing inspiration from the field of natural language processing [33], masked image modeling (MIM) techniques have emerged as highly effective learners in the vision domain [3, 25, 53]. This approach involves dividing images into small patches [14], with a high proportion of them being masked, and subsequently reconstructing them using a denoising autoencoder [47]. Notably, after the training phase, the decoder is discarded, leaving the encoder to serve as a feature extractor. MIM has been applied with success across a broad range of fields, and has had numerous extensions and improvements [45, 48, 2, 23, 18, 40, 19, 16, 17, 31, 9, 38].
@@
Siamese Masked Autoencoders. Building upon the work of masked autoencoders [25], Siamese Masked Autoencoders (SiamMAE) [23] have emerged as a new state-of-the-art in video propagation tasks such as video object segmentation [39], pose keypoint propagation [30], and semantic part propagation [56]. SiameseMAE uses a Siamese encoder [5] to process either pairs [23] or groups [31] of frames, randomly selected from a video. A key feature of SiameseMAE is its asymmetric masking technique: the initial frame undergoes no masking, thereby serving as a complete reference, while a substantial portion (up to 95%) of the second frame is masked. This setup encourages the network to accurately reconstruct the masked subsequent frames using the fully visible initial frame as a reference. The efficacy of SiameseMAE is believed to stem from its ability to effectively model object motion from videos and visual correspondence, learning the “propagation” and boundaries of objects from their observed positions in the past to their future locations, based on the few visible patches [23]. In this work, we show that explicit motion derived from videos is not mandatory for Siamese masked autoencoders to learn object-centric representations. Particularly, we demonstrate that the ability to recognize object boundaries and acquire propagation skills can be effectively learned from still images.
@@ 3 Method
We propose a novel self-supervised method, namely CropMAE, capable of learning valuable representations both from images and video frames. First, we create two augmented views ($V_{1}$ and $V_{2}$) of an input image ($I$) by randomly cropping, resizing and horizontally flipping the original image (Sec. 3.1). Second, we patchify [14] both views $V_{1}$ and $V_{2}$ (Sec. 3.2) and mask [33, 25] an extremely high portion of the second view ($V_{2}$) (Sec. 3.3). Both views are encoded in a Siamese [5] ViT encoder, with an additional positional embedding [14]. Third, a transformer [20] decoder reconstructs a target image $R$ (Sec. 3.4). The Siamese network and the decoder are trained by minimizing the L2 norm between the target $V_{2}$ and the reconstructed image $R$. After such pre-training, the decoder is discarded, and we use the encoder as a feature extractor on downstream tasks. With this setup, we demonstrate that meaningful data augmentations, in particular random crops, can provide meaningful information to pre-train visual encoders, injecting knowledge of motion and object boundaries. Figure 1 illustrates the main components of our method.
@@ 3.1 Cropping
  • Same Views. This setup corresponds to a direct adaptation of SiamMAE to images, in which the input image ($I$) is cropped once and serves both as the first view ($V_{1}$) for the key and value in the cross-attention decoder and as the second view ($V_{2}$) for the target for reconstruction. An illustration is given in Figure 2(a).
@@
  • Local-to-Global Views. In this setup, the unmasked view $V_{1}$ is a random cropped of the unmasked view $V_{2}$, and the masked view $V_{2}$ is another random cropped of the original image $I$. An illustration is provided in Figure 2(c).
  • Global-to-Local Views. Inversely, the masked view $V_{1}$ is a random cropped of the original image $I$, and the unmasked view $V_{2}$ is another random cropped of the unmasked view $V_{1}$. An illustration is provided in Figure 2(d).
@@ 3.3 Masking
Since both views are highly spatially redundant, a high masking ratio (above 75%) is usually necessary to create a challenging pretext task and to achieve optimized performances with masked autoencoders [25]. This is even more important in videos where both the spatial and temporal dimensions are highly redundant, requiring even higher masking ratios (90%) [45, 48, 17].
SiamMAE [23] employs a highly asymmetrical masking strategy, where the first frame is left completely visible while the second one is masked at $95\%$, which corresponds to $9$ visible patches out of the $196$ available when using a ViT/16 [14]. Using such a high masking ratio encourages the model to propagate the visible patches from the first frame to the second one and to learn temporal correspondences through motion [23]. However, employing a high masking ratio can make some examples ambiguous or may require additional knowledge beyond merely “propagating” patches from the unmasked view. For instance, if an object is only partially visible in the first view, while it is completely present (but masked) in the second one, the task becomes intractable if the model relies solely on the first view to reconstruct it. This prompts the model to learn a conceptual representation of the objects it encounters [25], enabling it to “hallucinate” what it partially sees when propagating past patches is either impossible or insufficiently informative.
@@ 3.4 Encoder and Decoder Architectures
Following [23], we use a Siamese ViT [14] encoder to process our two views and a vanilla Transformer [46] composed of cross-attention and self-attention layers as our decoder. Specifically, our decoder alternates between self-attention, where tokens of the masked image attend to each other, and cross-attention layers, where the tokens of the masked image attend to tokens of the visible image.
We train the Siamese architecture by minimizing the L2 loss between the normalized [25] pixel values of the view $V_{2}$ and the reconstruction $R$.
@@ 4.1 Experimental setup
Implementation details. Following previous methods [6, 23, 45], we use the ViT-S/16 as encoder architecture [14] for most of our experiments and fair comparisons with respect to other methods in the field. For the decoder, we employ a 4-layer Transformer [46] with a dimension $d_{\text{model}}=256$, where each block comprises a cross-attention layer, a feed-forward layer (of dimension $d_{\text{ff}}=2048$), and a self-attention layer. GELU activation functions [27] were utilized alongside a dropout rate of 10% [44]. We use the AdamW [35] optimizer and a base learning rate of $1.5e^{-4}$. The exhaustive list of hyper-parameters that we use can be found in the supplementary material.
Baselines. We compare our method with several state-of-the-art methods including MAE-ST [17], MAE [25], VideoMAE [45], and SiamMAE [23]. To the best of our knowledge, no official open-source code is available for SiamMAE, so we reimplemented it to compare the evolution of our performance during training, using the exact same hyperparameters described in the SiamMAE paper (Refer to the supplementary material). Our results are consistent with the ones reported in their paper [23]. However, we train for $400$ epochs instead of $2000$ to save computational resources.
Datasets. We pre-train our models on Kinetics-400 [32] (K400), on ImageNet [41] (IN), or on a subset of ImageNet (IN Subset). IN Subset contains $239,787$ randomly selected images, which corresponds to the number of videos in K400, for fair comparison between methods trained on K400 and ImageNet. During pre-training, we randomly sample an image (or a frame on K400), which is then processed following our methodology described in Section 3.
Downstream tasks. We evaluate our method on three propagation downstream tasks: video object segmentation (DAVIS-2017 [39]), human pose propagation (JHMDB [30]) and semantic part propagation (VIP [56]). These propagation tasks are framed as a semi-supervised problem, where the first annotated frame is provided, and the model is expected to propagate the segmentation mask to subsequent frames.
@@ 4.2 Results
We compare our method to previous works and present quantitative results in Table 1. We then provide some qualitative results of the reconstructed image and the downstream tasks respectively in Figures 3 and 4. The first part of the table displays results as reported in their original papers, under optimal training conditions in terms of both training duration and data volume. In the second part of the table, we report the outcomes achieved under our constrained training: either on K400 or on our ImageNet Subset, for a fixed duration of $400$ epochs.
Table 1: Comparison with prior work. We evaluate our method on three downstream tasks: video object segmentation (DAVIS-2017 [39]), human pose propagation (JHMDB [30]) and semantic part propagation (VIP [56]). Specifically, we compare our method with other methods trained on 400 epochs, on K400 [32] or on our ImageNet [12] Subset (IN Sub) for fair comparison. $\dagger$ refers to results reported in [23]. $\ddagger$ refers to our implementation, discussed in the supplementary materials.
@@
When trained for $2{,}000$ epochs on K400, SiamMAE achieves state-of-the-art performances on the three downstream tasks, and outperforms previous MAE methods such as MAE-ST [17], MAE [25] and VideoMAE [45]. However, considering a fixed budget of $400$ epochs, CropMAE achieves significantly better results than SiamMAE on DAVIS-2017 [39], both when trained on K400 and on our ImageNet Subset ($+0.7\%$ and $+2.5\%$ respectively). We believe that by explicitly transforming images through cropping, our pre-training method more quickly understands features useful for segmentation, such as object boundaries. On VIP [56], CropMAE still performs better than SiamMAE, although by a smaller margin ($+0.1$ when trained on ImageNet, and $+0.5$ when trained on K400). On JHMDB [30], CropMAE performs better than VideoMAE, but not the other methods.

### v2

@@ Abstract
Self-supervised pre-training of image encoders is omnipresent in the literature, particularly following the introduction of Masked autoencoders (MAE). Current efforts attempt to learn object-centric representations from motion in videos. In particular, SiamMAE recently introduced a Siamese network, training a shared-weight encoder from two frames of a video with a high asymmetric masking ratio (95%). In this work, we propose CropMAE, an alternative approach to the Siamese pre-training introduced by SiamMAE. Our method specifically differs by exclusively considering pairs of cropped images sourced from the same image but cropped differently, deviating from the conventional pairs of frames extracted from a video. CropMAE therefore alleviates the need for video datasets, while maintaining competitive performances and drastically reducing pre-training and learning time. Furthermore, we demonstrate that CropMAE learns similar object-centric representations without explicit motion, showing that current self-supervised learning methods do not learn such representations from explicit object motion, but rather thanks to the implicit image transformations that occur between the two views. Finally, CropMAE achieves the highest masking ratio to date (98.5%), enabling the reconstruction of images using only two visible patches. Our code is available at https://github.com/alexandre-eymael/CropMAE.
@@ 1 Introduction
Figure 1: CropMAE self-supervised pre-training. Given an input image ($V_{1}$), a second image is generated by performing a random crop and, optionally, a horizontal flip on the original image ($V_{2}$). We then patchify [14] both views and mask [33, 25] an extremely high portion of the second image (above 98.5%). Both views are encoded by a Siamese [5] ViT encoder, with added positional embedding [14]. A transformer [20] decoder reconstructs the masked image $R$ using self-attention layers on the tokens of the masked image and cross-attention layers between the tokens of the masked and unmasked images.
@@
Recently, Siamese Masked autoencoders (SiamMAE) [23] achieved state-of-the-art performance in numerous propagation tasks [56, 39, 30] by learning object-centric representations from videos. This method leverages a Siamese encoder [5] to process pairs of frames that are asymmetrically masked. Despite its impressive performance, SiamMAE faces two main limitations. Firstly, it is designed to only process video frames, not standalone images. Yet, image datasets are typically orders of magnitude larger than video datasets, and less computationally expensive to decode, making image-based pre-training more effective and scalable than video-based pre-training. Secondly, while SiamMAE reduces the need for the intense data augmentation found in contrastive learning methods, it still requires learning a conceptual understanding of the visual world, similar to most MIM techniques, thus requiring extensive training ($2{,}000$ epochs) on large datasets such as K400 [32] to reach state-of-the-art performances.
@@
In this work, we propose a novel self-supervised learning method, called CropMAE, that reframes the siamese-based paradigm introduced in SiamMAE in order to alleviate the need for video dataset, while keeping competitive performances on downstream tasks. Specifically, we use random views of the same image to simulate viewpoint changes, object transformations, motion, and occlusions. Our method can therefore leverage both image and video datasets, and train at a significantly faster pace than SiamMAE. Moreover, we demonstrate that CropMAE learns meaningful object-centric representations for downstream video tasks without explicit motion. Finally, unlike most MIM techniques, the pretext task of CropMAE is directly tractable based on the visible frame without the need to learn conceptual information about the world, which we believe is the reason for its faster training. An overview of our method is presented in Figure 1.
Contributions.
@@
We summarize our contributions as follows. (i) We introduce a novel pre-training method, CropMAE, based on sole images, which alleviates the need for video decoding and significantly accelerates training. The novel pretext task we introduce learns faster while quickly reaching good performances. (ii) We empirically demonstrate the feasibility of learning meaningful representations for downstream video tasks from still images or data distributions traditionally not associated with videos. Notably, this approach yields better results than training directly on video frames. (iii) We show, for the first time, that employing an extremely high masking ratio (98.5%, i.e., using only two visible patches for a ViT/16), surpassing those explored in existing studies, can be optimal and generate a meaningful and challenging self-supervised task.
@@ 2 Related Work
Visual self-supervised learning focuses on learning rich and generalizable representations of images or videos. This is typically achieved through pretext tasks [36, 37, 43, 8], enabling the learned representations to be applicable to a broad set of downstream tasks [15, 12, 56], either by fine-tuning the learned models for specific tasks, or by freezing the weights and training a linear classifier or an MLP on top of it. Key downstream tasks in the visual domain include image classification [8, 22, 6, 10, 52, 55, 3, 25, 19, 38], video classification [45, 48, 17, 19, 16, 38], object detection [22, 10, 25], and video segmentation [6, 10, 4, 23, 31]. Our method, CropMAE, is a new visual self-supervised representation learning method for propagation tasks [39, 56, 30].
Contrastive Self-Supervised Learning.
@@
Contrastive self-supervised learning [24] has been recognized as an effective method for feature extraction, applicable both to images [10, 22] and videos [42, 11]. This approach encourages the encoder to learn robust representations of the input data by minimizing the distance between representations of different augmented versions of the same image. Initially, it was common to enforce distinct images to have different representations in order to avoid representation collapse [13, 50, 8]. However, subsequent discoveries [26, 22] have shown that robust learning can be achieved even without imposing this constraint. Contrastive self-supervised learning has also been widely used for correspondence learning [49, 29], as it inherently learns to build representations that are invariant and robust to perturbations. Contrary to contrastive learning, CropMAE does not rely as extensively on data augmentations and is not subject to representation collapse issues.
Masked Image Modeling.
@@
Drawing inspiration from the field of natural language processing [33], masked image modeling (MIM) techniques have emerged as highly effective learners in the vision domain [3, 25, 53]. This approach involves dividing images into small patches [14], with a high proportion of them being masked, and subsequently reconstructing them using a denoising autoencoder [47]. Notably, after the training phase, the decoder is discarded, leaving the encoder to serve as a feature extractor. MIM has been applied with success across a broad range of fields, and has had numerous extensions and improvements [45, 48, 2, 23, 18, 40, 19, 16, 17, 31, 9, 38].
Siamese Masked Autoencoders.
@@
Building upon the work of masked autoencoders [25], Siamese Masked Autoencoders (SiamMAE) [23] have emerged as a new state-of-the-art in video propagation tasks such as video object segmentation [39], pose keypoint propagation [30], and semantic part propagation [56]. SiameseMAE uses a Siamese encoder [5] to process either pairs [23] or groups [31] of frames, randomly selected from a video. A key feature of SiameseMAE is its asymmetric masking technique: the initial frame undergoes no masking, thereby serving as a complete reference, while a substantial portion (up to 95%) of the second frame is masked. This setup encourages the network to accurately reconstruct the masked subsequent frames using the fully visible initial frame as a reference. The efficacy of SiameseMAE is believed to stem from its ability to effectively model object motion from videos and visual correspondence, learning the “propagation” and boundaries of objects from their observed positions in the past to their future locations, based on the few visible patches [23]. In this work, we show that explicit motion derived from videos is not mandatory for Siamese masked autoencoders to learn object-centric representations. Particularly, we demonstrate that the ability to recognize object boundaries and acquire propagation skills can be effectively learned from still images.
@@ 3 Method
We propose a novel self-supervised method, namely CropMAE, capable of learning valuable representations both from images and video frames. First, we create two augmented views ($V_{1}$ and $V_{2}$) of an input image ($I$) by randomly cropping, resizing and horizontally flipping the original image (Sec. 3.1). Second, we patchify [14] both views $V_{1}$ and $V_{2}$ (Sec. 3.2) and mask [33, 25] an extremely high portion of the second view ($V_{2}$) (Sec. 3.3). Both views are encoded in a Siamese [5] ViT encoder, with an additional positional embedding [14]. Third, a transformer [20] decoder reconstructs a target image $R$ (Sec. 3.4). The Siamese network and the decoder are trained by minimizing the L2 norm between the target $V_{2}$ and the reconstructed image $R$. After such pre-training, the decoder is discarded, and we use the encoder as a feature extractor on downstream tasks. With this setup, we demonstrate that meaningful data augmentations, particularly random crops, can generate rich and useful object-centric representations for propagation tasks without explicit motion.  Figure 1 illustrates the main components of our method.
@@ 3.1 Cropping
  • Same Views. This setup corresponds to a direct adaptation of SiamMAE to images, in which the input image $I$ is cropped once and serves both as $V_{1}$ and $V_{2}$. An illustration is given in Figure 2(a).
@@
  • Local-to-Global Views. In this setup, the masked view $V_{2}$ is a random crop of the original image $I$, and the unmasked view $V_{1}$ is another random crop of the masked view $V_{2}$. An illustration is provided in Figure 2(c).
  • Global-to-Local Views. Inversely, the unmasked view $V_{1}$ is a random crop of the original image $I$, and the masked view $V_{2}$ is another random crop of the unmasked view $V_{1}$. An illustration is provided in Figure 2(d).
@@ 3.3 Masking
Since both views are highly spatially redundant, a high masking ratio (above 75%) is usually necessary to create a challenging pretext task and to achieve optimized performances with masked autoencoders [25]. This is even more important in videos where both the spatial and temporal dimensions are highly redundant, requiring even higher masking ratios (90%) [45, 48, 17]. SiamMAE [23] employs a highly asymmetrical masking strategy, where the first frame is left completely visible while the second one is masked at $95\%$, which corresponds to $9$ visible patches out of the $196$ available when using a ViT/16 [14]. Using such a high masking ratio encourages the model to propagate the visible patches from the first frame to the second one and to learn temporal correspondences through motion [23]. However, employing a high masking ratio can make some examples ambiguous or may require additional knowledge beyond merely “propagating” patches from the unmasked view. For instance, if an object is only partially visible in the first view, while it is completely present (but masked) in the second one, the task becomes intractable if the model relies solely on the first view to reconstruct it. This prompts the model to learn a conceptual representation of the objects it encounters [25], enabling it to “hallucinate” what it partially sees when propagating past patches is either impossible or insufficiently informative.
@@ 3.4 Encoder and Decoder Architectures
Following [23], we use a Siamese ViT [14] encoder to process our two views and a vanilla Transformer [46] composed of cross-attention and self-attention layers as our decoder. Specifically, our decoder alternates between self-attention, where tokens of the masked image attend to each other, and cross-attention layers, where the tokens of the masked image attend to tokens of the visible image. We train the Siamese architecture by minimizing the L2 loss between the normalized [25] pixel values of the view $V_{2}$ and the reconstruction $R$.
@@ 4.1 Experimental setup
Implementation details.
Following previous methods [6, 23, 45], we use the ViT-S/16 as encoder architecture [14] for most of our experiments and fair comparisons with respect to other methods in the field. For the decoder, we employ a 4-layer Transformer [46] with a dimension $d_{\text{model}}=256$, where each block comprises a cross-attention layer, a feed-forward layer (of dimension $d_{\text{ff}}=2048$), and a self-attention layer. GELU activation functions [27] are utilized alongside a dropout rate of 10% [44]. We use the AdamW [35] optimizer and a base learning rate of $1.5e^{-4}$. The exhaustive list of hyper-parameters that we use can be found in the Appendix 6.1.
Baselines.
We compare our method with several state-of-the-art methods including MAE-ST [17], MAE [25], VideoMAE [45], and SiamMAE [23]. To the best of our knowledge, no official open-source code is available for SiamMAE, so we reimplemented it to compare the evolution of our performance during training, using the exact same hyperparameters described in the SiamMAE paper (refer to the supplementary material). Our results are consistent with the ones reported in their paper [23]. However, we train for $400$ epochs instead of $2000$ to save computational resources. Results for longer training can be found in the Appendix 6.2.
Datasets.
We pre-train our models on Kinetics-400 [32] (K400), on ImageNet [41] (IN), or on a subset of ImageNet (IN Subset). IN Subset contains $239,787$ randomly selected images, which corresponds to the number of videos in K400, for fair comparison between methods trained on K400 and ImageNet. During pre-training, we randomly sample an image (or a frame on K400), which is then processed following our methodology described in Section 3.
Downstream tasks.
We evaluate our method on three propagation downstream tasks: video object segmentation (DAVIS-2017 [39]), human pose propagation (JHMDB [30]) and semantic part propagation (VIP [56]). These propagation tasks are framed as a semi-supervised problem, where the first annotated frame is provided, and the model is expected to propagate the segmentation mask to subsequent frames.
@@ 4.2 Results
We compare our method to previous works and present quantitative results in Table 1. We then provide some qualitative results of the reconstructed image and the downstream tasks respectively in Figures 3 and 4. The first part of Table 1 displays results as reported in their original papers, under optimal training conditions in terms of both training duration and data volume. In the second part, we report the results achieved by our reproduced implementation of SiamMAE and CropMAE under our constrained training: either on K400 or on our ImageNet Subset, for a fixed duration of $400$ epochs, and for both ViT-S/16 and ViT-B/16.
Table 1: Comparison with prior work. We evaluate our method on three downstream tasks: video object segmentation (DAVIS-2017 [39]), human pose propagation (JHMDB [30]) and semantic part propagation (VIP [56]). Specifically, we compare our method with other methods trained on 400 epochs, on K400 [32] or on our ImageNet [12] Subset (IN Sub) for fair comparison. $\dagger$ refers to results reported in [23]. $\ddagger$ refers to our implementation.
@@
CropMAE (ours) ViT-B/16 IN Sub 400 60.9 57.9 63.8 32.8 44.3 72.3
When trained for $2{,}000$ epochs on K400, SiamMAE achieves state-of-the-art performances on the three downstream tasks, and outperforms previous MAE methods such as MAE-ST [17], MAE [25] and VideoMAE [45]. However, considering a fixed budget of $400$ epochs, CropMAE achieves significantly better results than SiamMAE on DAVIS-2017 [39], both when trained on K400 and on our ImageNet Subset ($+0.7\%$ and $+2.5\%$ respectively). We believe that by explicitly transforming images through cropping, our pre-training method more quickly understands features useful for segmentation, such as object boundaries. On VIP [56], CropMAE still performs better than SiamMAE, although by a smaller margin ($+0.1$ when trained on ImageNet, and $+0.5$ when trained on K400). On JHMDB [30], CropMAE only outperforms VideoMAE. We explain these inferior performances by noting that SiamMAE uses two different frames, resulting in complex human pose modifications, which likely helps the network understand human motion and perform better on JHMDB. Conversely, our random crops do not mimic these transformations. Yet, they help the network learn object boundaries more explicitly, making it more suited for segmentation tasks such as DAVIS.